mikewwe13:

Devonte is purchasing a blocked shaped like a right rectangular prism for a construction project. He knows the volume, V, and the area of the base, B, of the block. Devonte needs to be sure that the block is not too tall for his project. He wants to rearrange the formula V = Bh so that the height, h, of the block can be determined given the volume and the area of the base. Which equation should Devonte use ? A. h = B/V B. h = V/B C. h = V x B D. h = V - B

Vocaloid:

hint: take V = Bh and divide both sides by B

mikewwe13:

1

Vocaloid:

to divide by B, write a division sign, then write B.

Vocaloid:

for example, to divide V by B, write V, then write a division sign, then write B.

Vocaloid:

V divided by B = V/B now do the same thing with Bh

mikewwe13:

H divided by B = BH

Vocaloid:

divide Bh by B

Vocaloid:

Write down Bh write a division sign then write B then solve

mikewwe13:

Bh ÷ B = Bh

Vocaloid:

hint: B/B = 1 so we can cancel out Bh/B to get h

Vocaloid:

so h = V/b is your answer (answer choice B)

mikewwe13:

The mapping shows the relationship between the variables x and y of function f, where the domain is set x and the range is set y. Which equations are true? Select the two correct equations. A. f(-3) = 21 B. f(9. 4) = 5 C. f(2) · 8 = 16 D. f(0) - 5 = 16 E. f(6.8) - 6.8 = 0

Vocaloid:

is there a picture that goes with this?

mikewwe13:

1 attachment
Vocaloid:

hint: when x = -3, what does f(x) = ?

mikewwe13:

-3

Vocaloid:

the arrow tells you what f(x) equals to look at the arrow at x = -3 where is it pointing? what number?

mikewwe13:

21

Vocaloid:

good, so f(-3) = 21 making A true (we still need to keep going)

Vocaloid:

B. f(9. 4) = 5 this implies that x = 9.4 and f(x) = 5 (this is not true because there is no x value 9.4) let's keep going

Vocaloid:

look at the pic again what is f(2)?

mikewwe13:

f(2)

Vocaloid:

f(2) means the y-value when x = 2 look at the pic what is the arrow at x = 2 pointing at?

mikewwe13:

16

Vocaloid:

good, this means f(2) = 16 C states that f(2) * 8 = 16 which is clearly not true since 16*8 is not 16 so let's keep moving

Vocaloid:

D. f(0) - 5 = 16 using the same logic from before, what is f(0)? answer this by looking at the pic, where x = 0, to find the matching f(x) value

mikewwe13:

x = 0

Vocaloid:

ok, but what is f(0)? see where the arrow is pointing?

mikewwe13:

the answer is 0

Vocaloid:

look at the arrow where x = 0. where is it pointing?

mikewwe13:

21

Vocaloid:

good so f(0) = 21 making f(0) - 5 = 21 - 5 = 16 (true!) so our answers are A & D

Vocaloid:

can we open a new question please